Let F(x, y, z)=(x + 2xy, y – 3zy, z + x2) Find the divergence and curl of F.
Answer #1

F(49,7) = <x+244, Y-3zy, 2+xz> =< Ft, F2, F3 > Fi = X+2x4, F2 = y-374, F3 =Z+XZ ofs + dfz Divergence of F- = + gp Ze up he JA